OpenStudy (anonymous):

Compute and write your answer in scientific notation. 6.3 x 10-7 3.25 x 10-12

OpenStudy (anonymous):

Try this: (6.3*3.25) then do the powers of ten

OpenStudy (bloomlocke367):

are you dividing or multiplying this one?

OpenStudy (anonymous):

dividing

OpenStudy (anonymous):

Ok same thing

OpenStudy (bloomlocke367):

okay... just do what you did in the last one...

OpenStudy (anonymous):

That just makes it easier

OpenStudy (bloomlocke367):

6.3/3.25=?

OpenStudy (anonymous):

6.3/3.25=1.938 10-7/10-12=105 1.938*105

OpenStudy (bloomlocke367):

\(\Huge\color{teal}{CORRECT!!!}\)

OpenStudy (anonymous):

awesome thank you!!:]

OpenStudy (bloomlocke367):

you're welcome :)

OpenStudy (anonymous):

thanks for all you help bloomlocke387 have a goodnight!

OpenStudy (bloomlocke367):

I wrote you a testimonial.. did you get it? and of course :) that's what I'm here for

OpenStudy (anonymous):

I did thanks again!:] night!

OpenStudy (bloomlocke367):

gn :)